Please help me out with this question:(explain)


Q)How many electrons weigh 1kg ?  (IIT 2002)


Ans)10^31/9.108

Grade:11

1 Answers

Pramod J AskiitiansExpert-IIT-B
12 years ago

Dear student,

mass of an electron = 9.10938188 × 10-31 kilograms

number of electrons to weigh 1 kg = 1/9.10938188 × 10-31 

= 10^31/9.109

i hope it explains !!